numerical question How many grams of ammonium iodide are in 355 ml of 0.995 M NHal solution?
Answer #1

volume , V = 355 mL

= 0.355 L

use:

number of mol,

n = Molarity * Volume

= 0.995*0.355

= 0.3532 mol

Molar mass of NH4I,

MM = 1*MM(N) + 4*MM(H) + 1*MM(I)

= 1*14.01 + 4*1.008 + 1*126.9

= 144.942 g/mol

use:

mass of NH4I,

m = number of mol * molar mass

= 0.3532 mol * 1.449*10^2 g/mol

= 51.2 g

Answer: 51.2 g